OpenStudy (anonymous):

what are the points for this system of equalities?: x+y is greater than or equal to 2+4, y<4 ? i

OpenStudy (anonymous):

to graph \(x+y \ge 6\) using intercept form of a line...\[\frac{x}{a}+\frac{y}{b}=1\]so the line will have x = y = 6 as intercepts..\[\frac{x}{6}+\frac{y}{6}=1\] 

OpenStudy (anonymous):

... the points in the over-lap region are the points for the given system of inequalities....

OpenStudy (anonymous):

if y<4, x>2...

OpenStudy (anonymous):

... the shaded region or over-lap region indicates that while y value is less than 4, the x value is greater than 2, maintaining the equality given in the first that \[x+y \ge 6\]

OpenStudy (anonymous):

... got the idea @felgail ?